A rectangular open box is have a square base, and its volume is to be 125 inches^3. The cost Per in^2 Of the material for the bottom Is 8$ and for the sides is 4$.


A. Find the mathematical model Expressing The total cost Of the material As a function of the edge Lenght Of the square base.


B. What is the total cost If the square Base has edge Lenght of 4 inches


1
Expert's answer
2020-10-03T15:42:42-0400

A. Denote length of the side of the square base by "x." The volume "V" of a rectangular box is equal to "x^2h", where "h" is the heigth of the box. Therefore, "h=\frac{V}{x^2}=\frac{125}{x^2}". The area of the lateral sides is equal to "4\frac{125}{x^2}x=\frac{500}x". Thus, the mathematical model expressing the total cost of the material can be written in the following form:

"m(x)=8x^2+4\frac{500}{x}=8x^2+\frac{2000}{x}"


B. If the square base has edge length of "x=4" inches, then its total cost is


"m(4)=8\cdot 4^2+\frac{2000}{4}=128+500=628."